solve quadratic equation by completing the square
-x^2-2x=5

Answer by mathstutor494(120)

-x^2-2x=5
> x^2+2x+5=0
> x^2+2x+1+4=0
> (x+1)^2=-4
>(x+1)^2=4i^2
So x+1 = +-2i
x= -1+-2i